Consider the titration of 60.0 mL of 0.0400 M HONH2 (a weak base; Kb = 1.10e-08) with 0.100 M HIO4. Calculate the pH after the following volumes of titrant have been added:


(a) 0.0 mL

pH =




(b) 6.0 mL

pH =




(c) 12.0 mL

pH =




(d) 18.0 mL

pH =




(e) 24.0 mL

pH =




(f) 40.8 mL

pH =
Answer #1

Aug) 3 оfurn * * ginen: Molarity of 1104 = o. 100M volume 9 N4201 - 6000m L=60.0x16L Molasity g MH204 = 0.0400M (a) volume ofThus, Cou)-*-.2.09707x105 polo -log Con) = -log(2.09707x165) = 4.67838707 PH = 14 - pon = 14 - 4.67838707 = 9.32161293 = 9.Applying the Henderson - Hasselbalch equallon pol = pkb + log (But to [B] whose [but] - [MHz out [B]= [NHon] Pk6= -log(Kb) =Substituting the values in ean POH = 7.95860731 +log (0.0090909099 0.0272727273 = 7.481 48605 PH= 14- poH = 14 - 7.48148605 =(d) Volume g 1104 - 18.0ml = 18.0x103L Moles of Hoq = 0.100 mal x 18.8x10²4 = 0.0018 mol Constructing ICE table for the reactPon= 8.4357 2857 PHI 11 - POW - 14 - . 435+18 51 = 5.56427143 = 5.56 (e) volume of 1104 = 24 mL = 24x16L Moles g 1104 = 0.100Constructing ICE table for the reaction: SW2OH (aq) + Học (49) » N, OH + H20(4) 0.0024 Inittal 0.0024 Cmoles) Change -0.0024[amongst = Moles of NH 2 O Hat volume g 11047 Volume oj NIKOM = 0.0024 mol = 0.0024 moll 24+60) xio3L 84x18L = 0.0285714286MM (0.02.85714286-w) 9.09090909 x107 = 2.9994026 X108 - 9.09090909xione my?+ 9.09090909 xiołne - 2.5974026x108-0 Solveig for aTatal volume Volume g 1104+ Volume oj Nyon (40.8+60) mL = 100.8x10?L (M304] - Excasse moles g 4104 Total volume 0.00 168 mol